OpenStudy (anonymous):

Which statement about the function below is true? f(x) = x^3 5x^2 25x + 125 = (x + 5)(x 5)(x 5) (Points : 4) The root 5 has a multiplicity of 1. The root 5 has a multiplicity of 2. The root 5 has a multiplicity of 1. The root 5 has a multiplicity of 3.

OpenStudy (anonymous):

@undeadknight26 @InfiniteMinds

OpenStudy (anonymous):

the first 2 are minus and so are the last two

OpenStudy (anonymous):

im confused o-o

OpenStudy (anonymous):

@andrewhaze

OpenStudy (e.mccormick):

So you are saying: \(f(x) = x^3-5x^2-25x + 125 = (x + 5)(x-5)(x-5)\) And basically what they want to know is when would x=5.

OpenStudy (anonymous):

yes

OpenStudy (andrewhaze):

ill let @e.mccormick help xP

OpenStudy (e.mccormick):

OK. So, for any \((x\pm a)\) factor you flip the sign to say \(x=\mp a\) So the factor (x+1) means x = -1. Now, if you apply that, how many are equal to 5?

OpenStudy (anonymous):

2

OpenStudy (e.mccormick):

Yep.

OpenStudy (anonymous):

so B

OpenStudy (e.mccormick):

And there you go. Graphically you can see it just touches, not crosses, at that point: https://www.desmos.com/calculator/g9tddezcnm

OpenStudy (anonymous):

Thank youuu!!!